OpenStudy (anonymous):

Given f(x) and g(x) = f(x) + k, look at the graph below and determine the value of k

OpenStudy (anonymous):

Can you provide the graph?

OpenStudy (anonymous):

heres the graph

OpenStudy (anonymous):

so g(x) = -1/3x+1 f(x) =-1/3x -3 if g(x) = f(x)+ k -1/3x+1 = -1/3x-3+k 1=-3+k k=4

OpenStudy (anonymous):

count

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome

OpenStudy (aum):

g(x) = f(x) + k Adding a constant "k" to f(x) has the effect of shifting the graph f(x) UP by "k" units. Look at the y-intercept of f(x) and look at the y-intercept of g(x). By how many units has the graph of f(x) moved up to form g(x)? By 4 units. So k = 4.